(a) Calculate the wavelengths in air at 20°C for sounds in the maximum range of human hearing, 20 Hz to 20,000 Hz. (b) What is the wavelength of an 18-MHz ultrasonic wave?


1
Expert's answer
2021-03-21T11:36:23-0400

(a) The speed of sound at 20°C equals "343\ m/s". Then, we can find the wavelengths from the wave speed formula:


"v=f\lambda,""\lambda_{20\ Hz}=\dfrac{v}{f}=\dfrac{343\ \dfrac{m}{s}}{20\ Hz}=17.15\ m,""\lambda_{20\ kHz}=\dfrac{v}{f}=\dfrac{343\ \dfrac{m}{s}}{20\cdot10^3\ Hz}=1.7\cdot10^{-2}\ m"

(b)

"\lambda=\dfrac{v}{f}=\dfrac{343\ \dfrac{m}{s}}{18\cdot10^{6}\ Hz}=1.9\cdot10^{-5}\ m."